the length of a rectangle is 5m less than twice its width. If its width is increased by 5m and its length by 2m, the new area would be 81 sq m more than the original. Find the dimensions of the rectangle.
Answer by edjones(8007)

L=2w-5
A=w(2w-5)
(w+5)((2w-5)+2)=81+w(2w-5)
(w+5)(2w-3)=81+2w^2-5w
2w^2+7w-15=2w^2-5w+81
12w=96
w=8
L=16-5=11
I checked this and its ok.
Ed